OpenStudy (bbb911):

Amy paid $63.75 for a pair of walking shoes during a 15%-off sale. What was the regular price?

OpenStudy (bbb911):

wouldn't the 15 percent be 0.15?

OpenStudy (bbb911):

cos that's what i was taught that you divide the percents by a hundred

OpenStudy (anonymous):

Yes. You add the 0.15 (discounted) to 1(original) and get 1.15. You then multiply it by the sale price to get the original price.

OpenStudy (bbb911):

what but what is the original dont i add it 0.15 to the sale price tho to get the original price? or cos im kind of confused sorry

OpenStudy (johnweldon1993):

Original price * (1 - percentage sale )= current price original price = (current price)/(1 - percentage sale ) original price = x current price = 63.75 percentage sale =15% x = 63.75/(1 - .15) x= 63.75/.85 x= 75 So the original price comes out to 75 dollars

OpenStudy (johnweldon1993):

Hope that helps clarify anything :)

OpenStudy (bbb911):

hopefully if i can't figure it out if i can't is it possible to help me ?

OpenStudy (johnweldon1993):

Of course :)

OpenStudy (bbb911):

kk thnk u :)

OpenStudy (johnweldon1993):

anytime! :)

OpenStudy (bbb911):

Original price * (1 - percentage sale )= current price when you put this the 1 - the percentage sale or

OpenStudy (johnweldon1993):

I'm not too sure what you're asking there.... What you have quoted above...is how you fine the current price... The current price...would have been found by multiplying the original price by the sale price right? Well..we know the current price...and we want to find out how much it was originally...so we take the current price....and we DIVIDE it by the opposite of how much it was on sale for... a.k.a it was on SALE for 15 percent....so there was 85 percent left right/ so \[Original = \frac{ Current }{ 1 - percentage }\] Your percentage was 15%..or .15 so \[Original = \frac{ Current }{ 1 - .15 }\] With the current price entered and 1 - .15 equaling .85 we have \[Original = \frac{ 63.75 }{ .85 }\] Did that kind of help?

OpenStudy (bbb911):

ohhhh! i actually did tht but wasn't sure i got the .85. just wan't sure if was right . sorry bout that tahnk :) let me figure it out now :)

OpenStudy (johnweldon1993):

Yes you were correct :) very good! ...and okay let me know how it goes :)

OpenStudy (bbb911):

kk will do thnks :)

OpenStudy (bbb911):

Got it 75 dollars, and its correct also on my book :) thank you for the method you sir are the best :) thanks again :)

OpenStudy (johnweldon1993):

Very good! and don't worry about it...anytime :)